# Detecting Spatial Dependence in Transcriptomics Data using Vectorised Persistence Diagrams

### v1

@@ 3.2 Summarising Persistence Diagrams
Persistence diagrams consist of multi-sets of birth- and death pairs and are thus challenging to study from a statistical perspective. In fact, there exists no natural way to directly understand them in as a Hilbert space that allows for the application of standard statistical tools. Indeed, inherent distances between diagrams such as the Bottelneck or Wasserstein distances are expensive to compute and do not induce a norm on the space of persistence diagrams. Potential approaches for fitting PH into a statistical setting is the use of summary statistics, as for example explored by Byers et al. [11] in the context of spatial dependency detection. Moreover, to retain the multiscale properties of PH, functional summaries can be employed to understand PDs in a Hilbert space setting.
@@ 3.2.1 Summary Statistics
While various one-number descriptors, such as persistence entropy or lifetime-derived statistics are available, we choose to include total lifetime (TL) into our study based on the findings of Byers et al. [11]. In particular, the total lifetime is computed as
$\operatorname{\mathrm{TL}}=\sum_{i=1}^{m}q_{i}-p_{i}$ (3.2)
i.e. the sum of the lifetimes of persistence pairs, defined as the difference between the death and birth scales of topological features.
@@ 3.2.2 Functional Summaries
Our analysis will focus on using functional summaries that map PDs to (piecewise) continuous functions from $\mathds{R}\to V$ where $V$ is a suitable vector space [4]. Note that in related literature these types of functional summaries are often called persistence curves. In particular, we compare Betti curves and persistence landscapes, which can be conveniently compared via suitable norms between these functions. We hypothesise that these functional summaries retain more complete information from the whole PD than single number summaries (as e.g. studied by Byers et al. [11]) while allowing for the application of statistical methods in a theoretically well-founded manner [9, 34].
@@
Betti curves are defined as $\operatorname{\mathrm{b}}:\mathds{R}\to\mathbb{N}$ where
$\operatorname{\mathrm{b}}(\delta)=|\{(p_{i},q_{i})\in\operatorname{\mathrm{PD}}\mid\delta\in[p_{i},q_{i}]\}|.$ (3.3)
That is, Betti curves sum up the number of topological features from the persistence diagram that persist at each scale of the filtration. Given one Betti curve, b, we can then take its $L^{p}$ norm as a natural descriptor:
@@
This $L^{1}$ norm then is closely related to computing total lifetime in terms of the information it summarises [34]. To compare multiple Betti curves, $\operatorname{\mathrm{b}}_{1}$ and $\operatorname{\mathrm{b}}_{2}$, we naturally extend this notion and use the $L^{p}$ distance between Betti curves, that is
@@
A frequently used tool for statistical analyses are persistence landscapes, which have been proven to follow a central limit theorem, even enabling the application of standard parametric tools [9]. First, the rank functions are defined for each persistence pair given as
$\lambda_{(p_{i},q_{i})}(\delta)=\begin{cases}\delta-p_{i}&\text{ if }p_{i}<\delta\leq\frac{p_{i}+q_{i}}{2}\\ -\delta+q_{i}&\text{ if }\frac{p_{i}+q_{i}}{2}<\delta<q_{i}\\ 0&\text{ otherwise}\\ \end{cases}$ (3.6)
@@
$\Lambda(k,\delta)=\text{kmax}\{\lambda_{(p_{i},q_{i})}(\delta)\}_{i\in I}$ (3.7)
where kmax is the function giving the k-th largest value of a set. The persistence landscape then is the collection of these functions. Further, we can define the $L^{p}$-norm for persistence landscapes as:
@@ 4.1 Computing Persistent Homology from Spatial Omics Data
Now, to develop our spatially variable gene detection method, we define a spatial graphs for omics data in the manner that best matches the underlying sequencing technology [39]. Amongst these technologies, spatial measurements are collected across varying resolutions. For example, for data with continuous spatial coordinates, such as high resolution measurements of single cells, we compute the Delaunay triangulation to find an appropriate triangular covering of the space. In comparison, Visium, one of the most popular sequencing methods, measures spots on a hexagonal grid with spots having diameters of $55\mu m$ and centers being a fixed distance of $100\mu m$ apart [26]. Motivated by this, we choose a spatial graph connecting each grid point to its six directly adjacent neighbouring spots for further analysis. For both of these graph choices we then apply simplicial-complex based filtrations. Lastly, for data measured on rectangular grids, we connect each spot to its four adjacent neighbours and apply cubical persistence to capture the rectangular nature of these grid cells.
Throughout this work we focus on computing zero-dimensional persistent homology, which allows us to investigate the spatial patterns given by connected components. In particular, we are interested in encoding peaks in spatial gene expression and therefore construct super-levelset filtrations using gene expression as vertex feature as described in Section 3.1. Then, the resulting persistence diagrams encode key information on clustering behaviour. Here, we choose to use a super-levelset approach rather than a sub-levelset filtration because we reason that spatial maxima will be more suitable for revealing spatial trends than spatial minima for these data. This choice is informed by the fact that omics data is affected by technical noise and sparsity. Indeed, dropout is a well known measurement error describing the common phenomena of not observing a transcript even though it is present. In practice, this effect leads to a large proportion of excess zeros in gene expression values [20], which could negatively affect the interpretation of local minima for our purpose. We therefore argue that spatial peaks will be more robust to such sparsity.
For each gene, whose spatial pattern we want to assess, we then compute persistence diagrams and persistence summaries as detailed in Section 3.1. By convention, we bound death value of each persistence pair by minimum filtration value. That is, rather than allowing one persistence feature to achieve an infinite death value, we bound it by the lowest gene expression value observed in the data. Note that our construction is similar to Byers et al. [11], but we further extend their methods to also consider cubical persistence for rectangular grids and investigate the use of functional summaries for spatial dependence testing as detailed below.
@@ 4.2 One-Sample Permutation Testing for Spatial Dependence
  • $H_{0}$: Expression values are randomly distributed in space
  • $H_{1}$: Expression values are not spatially random
We thus retain the original gene expression values as well as the spatial graph and randomly permute which expression value is assigned to each location to simulate spatial randomness, that is the lack of any spatial pattern. Then, significance is assessed as follows:
@@
where $\mathds{1}$ is the indicator function summing up the number of times the $L^{p}$-norm, $||\cdot||_{p}$, between the observed Betti curve $\operatorname{\mathrm{b}}_{\text{obs}}$ and the empirical mean Betti curve under the null model $\widebar{\operatorname{\mathrm{b}}_{0}}$ is more extreme than the difference between each individual Betti curve under the null model and their overall mean.
Note that we compute $\widebar{\operatorname{\mathrm{b}}_{0}}$, the mean Betti curve given the null hypothesis, as the mean across all $n_{\text{perm}}$ permutations as well as the observed Betti curve while also adding 1 to both the nominator and denominator. This bias correction is included to avoid zero p-values and bias the test towards more conservative estimates by including this pseudo-count [31]. We thus conduct a two-sided, one-sample hypothesis test for each spatial feature. Whenever multiple genes are compared simultaneously, we further recommend to apply a suitable multiple testing adjustment such as Benjamini-Hochberg procedure to control the false discovery rate [7], which we will apply throughout our experiments. Across our experiments, the number of permutations is set to 1000 giving a practical compromise with regards to computational efficiency considerations.
Beyond statistical reasoning, there also exist more deterministic approaches to spatially variable gene discovery. Here, the goal is to either rank features by the strength of their spatial dependence, or to find a predefined number of the top most spatially variable features. In practice, these approaches are key for both explanatory analysis and for feature selection as a first step to uncover and understand spatial patterns in data [6, 13]. For further analysis, we thus choose to rank spatially expressed features by their empirical p-values following similar statistical approaches [24].
@@ 4.3 Alternative Spatial Variability Detection Methods
In order to evaluate the performance of our spatial testing approach we compare our methods with three standard methods for detecting spatial dependence.
@@
Moran’s I is one of the oldest and most widely-used measures for spatial auto-correlation. It is known to be a suitable baseline for SVG benchmarking [24] and can be computed as
$I=\frac{n_{obs}}{\sum_{i,j}w_{i,j}}\frac{\sum_{i=1}^{n_{obs}}\sum_{j=1}^{n_{obs}}w_{i,j}(f_{i}-\bar{f})(f_{j}-\bar{f})}{\sum_{i=1}^{n_{obs}}(f_{i}-\bar{f})^{2}}.$ (4.2)
We compute it using the spatial adjacency matrix to define the weights $w_{i,j}$. The value of Moran’s I is directly interpretable, ranging from $-1$ indicating negative autocorrelation, to 00 corresponding to spatial randomness, and $1$ indicating positive autocorrelation. Further, for statistical analysis we use Moran’s I in the same randomised permutation testing framework as the previously introduced topological summaries explained in Section 4.2.
@@
sepal simulates diffusion processes, which models a stochastic process that tracks the spread of a feature on the spatial domain over time as it tends towards randomness. Diffusion is simulated on either a rectangular or hexagonal grid and sepal then measures diffusion times, which are used to assess spatial patterns and measure how long it takes the diffusion process to converge. The higher the diffusion time the stronger the spatial signal. Note that this method does not conduct any statistical testing and thus does not provide p-values. Instead, it provides a ranking of genes by their diffusion times. Moreover, sepal is only applicable to hexagonal or rectangular grids. Hence, whenever data is given with unstructured coordinates, we follow the implementation by Andersson and Lundeberg [6] to convert spatial measurements to a rectangular grid.
@@
Finally, SpatialDE applies spatial covariance testing via Gaussian process regression, an class of models originating from geostatistics. In a nutshell, gene expression is modelled via a multivariate normal model that includes both a non-spatial and a spatial covariance term. The model fitted with spatial covariance is then compared to a null model of spatial independence without this spatial component. Likelihood testing is used and p-values are computed analytically using the $\chi^{2}$ distribution with one degree of freedom. As a multiple testing correction, we follow Svensson et al. [38] and report q-values, which specifically control for the positive false discovery rate [37]. SpatialDE assumes normally distributed residual noise, we follow the preprocessing stpdf by Svensson et al. [38]. First, Anscombe’s transformation, a variance stabilising method for negative binomial data, is applied. Second, log total count values are regressed out to achieve independence from the total counts per spatial location.
@@ 4.4 Datasets and Preprocessing
We use three real transcriptomics datasets for our investigation. First, we analyse spatial data of a mouse olfactory bulb (referred to as MOB, replicate 1) and a breast cancer tumour sample (breast, replicate 1) from Ståhl et al. [36], standard datasets commonly used to test spatial variability detection methods. Both have been sequenced using 1k spatial transcriptomics arrays, which measure spots on rectangular grids. Each spot has a diameter of $100\mu m$ and encompasses 10-100 cells. Further, we study one sample from Kuppe et al. [23] (sample IZP10), which consists of spatial measurements of spots on a hexagonal grid. This sample has been taken from the heart of a patient that suffered from a heart attack from tissues with necrotic areas.
@@
Prior to analysis, peprocessing and quality control of omics data are crucial stpdf for accounting for technical errors as well as filtering out noise. Indeed, normalisation is a key step in spatial omics data analysis as many downstream tasks, including the detection of spatially variable features, rely on adequate preprocessing choices to account for technical effects [20]. Following best practices [20] and alternative spatial variability studies [16], we decide on the following pre-processing strategy and select all genes that have a minimum total expression count of $10$, and are present in at least 1% of locations. Further, we remove all cells with fewer than $10$ counts. These stpdf are done to avoid analysing cells or gene for which hardly any biological signals have been detected. Finally, we exclude all ribosomal and mitochondrial genes from analysis as these indicate technical artifacts and are not related to the mRNA, i.e. protein-coding genes that are the focus of transcriptomics data analysis [28]. For each real dataset, we then identify the $500$ most variable genes using $\texttt{seurat}_{\texttt{v3}}$ for further analysis. That is, we select a set of genes that vary most notably across cells. Our goal will then be to detect whether these genes that differ noticeably between dislocated cells also show distinct spatial patterns. Note that for the simulated data we skip these quality control stpdf as we want to compare all SVG detection methods across all simulated features.
@@
Further, as conventionally done in spatial omics data analysis, we decide to pre-process the expression values i.e. count data in the manner best fitting the SVG detection method [24, 13]. For SpatialDE we apply the normalisation strategy preferred for this method as detailed in Section 4.3 [38]. For sepal, we apply the recommended shifted logarithm normalisation by applying the transformation $\log(f+2)$ to each feature $f$. This transformation is applied to stabilize the variance across the dataset to counteract variable sampling effects such as overdispersion and reduce the skewedness of the feature’s distribution. The pseudo-count of $2$ is added to enable the log-transformation also for zero-values counts. To ensure comparability across spatial variability tests, we similarly choose to compute our persistence-based test as well as Moran’s I from the same $\log(f+2)$-transformed features. Note however, that this pre-processing strategy is not the only viable choice and our method can be flexibly applied across varying pre-processing choices [2]. In general, given that our persistence-based methods directly scales through each gene’s expression values, we highly recommend to apply our framework to transformed and normalised features, but we leave the choice of the best data-dependent transformation to further investigation [2].

### v2

@@ 3.3 Summarising Persistence Diagrams
Persistence diagrams consist of multi-sets of birth and death pairs and are thus challenging to study from a statistical perspective. In fact, there exists no natural way to directly understand them in a Hilbert space that allows for the application of standard statistical tools. Moreover, inherent distances between diagrams such as the Bottleneck or Wasserstein distances are expensive to compute. Thus, to gainfully use PH in a statistical setting, we can use summary statistics, as for example explored by Byers et al. [11] in the context of spatial dependency detection. Moreover, to retain the multiscale properties of PH, functional summaries can be employed to understand persistence diagrams in a Hilbert space setting. Fig. 1.4. illustrates the examples of three of such persistence summaries, which we will now explain in more detail.
@@ 3.3.1 Summary Statistics
While various one-number descriptors such as persistence entropy or lifetime-derived statistics are available, we choose to include total lifetime (TL) into our study based on the findings of Byers et al. [11]. In particular, the total lifetime is computed as:
$\operatorname{\mathrm{TL}}=\sum_{i=1}^{m}(q_{i}-p_{i}).$ (4)
That is, total lifetime is the sum of the lifetimes of persistence pairs, defined as the difference between the death $q_{i}$ and birth $p_{i}$ scales of topological features $(p_{i},q_{i})\in\operatorname{\mathcal{D}}$.
@@ 3.3.2 Functional Summaries
Our analysis will focus on using functional summaries that map persistence diagrams to (piecewise) continuous functions from $\mathds{R}\to V$ where $V$ is a suitable vector space [4]. Note that in related literature these types of functional summaries are often called persistence curves. In particular, we compare Betti curves and persistence landscapes, which can be conveniently compared via suitable norms between these functions. We hypothesise that these functional summaries retain more complete information from the whole persistence diagram than single number summaries (as e.g. studied by Byers et al. [11]) while allowing for the application of statistical methods in a theoretically well-founded manner [9, 37].
@@
Betti curves are defined as $\operatorname{\mathrm{b}}\colon\mathds{R}\to\mathbb{N}$ where
$\operatorname{\mathrm{b}}(\delta)=|\{(p_{i},q_{i})\in\operatorname{\mathcal{D}}\mid\delta\in[p_{i},q_{i}]\}|.$ (5)
That is, Betti curves sum up the number of topological features from the persistence diagram that persist at each scale of the filtration. Specifically, we filter from the highest vertex feature value, $\max(f_{i})$ to $\min(f_{i})$. Given one Betti curve, $b$, we can then take its $L^{p}$ norm as a natural descriptor:
@@
This $L^{p}$ norm then is closely related to computing total lifetime in terms of the information it summarises [37]. To compare multiple Betti curves, $\operatorname{\mathrm{b}}_{1}$ and $\operatorname{\mathrm{b}}_{2}$, we naturally extend this notion and use the $L^{p}$ distance between Betti curves, that is
@@
A frequently used tool for statistical analyses are persistence landscapes, which have been proven to follow a central limit theorem, even enabling the application of standard parametric tools [9]. First, the rank functions are defined for each persistence pair, $(p_{i},q_{i})\in\operatorname{\mathcal{D}}$, given as
$\lambda_{(p_{i},q_{i})}(\delta)=\begin{cases}\delta-p_{i}&\text{ if }p_{i}<\delta\leq\frac{p_{i}+q_{i}}{2}\\ -\delta+q_{i}&\text{ if }\frac{p_{i}+q_{i}}{2}<\delta<q_{i}\\ 0&\text{ otherwise}\\ \end{cases}$ (8)
@@
$\Lambda(k,\delta)=\text{kmax}\{\lambda_{(p_{i},q_{i})}(\delta)\}_{i\in I}$ (9)
where kmax is the function giving the k-th largest value of a set. The persistence landscape then is the collection of these functions. Further, we can define the $L^{p}$ norm for persistence landscapes as:
@@
where the vertex feature $f_{i}$ associated with $x_{i}\in X$ determines the minimum and maximum value of our filtration parameter $\delta$.
@@ 4.1 Computing Persistent Homology from Spatial Omics Data
Now, to develop our spatially variable gene detection method, we define spatial graphs from omics data in the manner that best matches the underlying sequencing technology [42]. Amongst these technologies, spatial measurements are collected across varying spatial resolutions. For example, for data with continuous spatial coordinates, such as high resolution measurements of single cells, we compute the Delaunay triangulation to find an appropriate triangular covering of the space [29]. In comparison, Visium, one of the most popular sequencing methods, measures spots on a hexagonal grid with spots having diameters of $55\mu m$ and centres being a fixed distance of $100\mu m$ apart [27]. Motivated by this, we choose a spatial graph connecting each grid point to its six directly adjacent neighbouring spots for further analysis. For both of these graph choices we then apply simplicial-complex based filtrations as explained in Section 3. Lastly, for data measured on rectangular grids, we connect each spot to its four adjacent neighbours and apply cubical persistence [46] as referenced in Section 3 to capture the rectangular nature of these grid cells.
Throughout this work we focus on computing zero-dimensional persistent homology, which allows us to investigate the spatial patterns given by connected components. In particular, we are interested in encoding peaks in spatial gene expression and therefore construct superlevel-set filtrations using gene expression as the vertex feature for constructing the filtrations as described in Section 3. Then, the resulting persistence diagrams encode key information on spatial clustering behaviour of gene expression values across the tissue sample. Here, we choose to use a superlevel-set approach rather than a sublevel-set filtration because we reason that spatial maxima will be more suitable for revealing spatial trends than spatial minima for these data. This choice is informed by the fact that omics data is affected by technical noise and sparsity. Indeed, dropout is a well-known measurement error describing the common phenomena of not observing a transcript even though it is present. In practice, this effect leads to a large proportion of excess zeroes in gene expression values [20], which could negatively affect the interpretation of local minima for our purpose. We therefore argue that spatial peaks will be more robust to such sparsity.
For each gene whose spatial pattern we want to assess, we then compute persistence diagrams and persistence summaries as detailed in Section 3. By convention, we bound the death value of each persistence pair by the minimum filtration value. That is, rather than allowing one persistence feature to achieve an infinite death value, we bound it by the lowest gene expression value observed in the data. Note that our construction is similar to Byers et al. [11], but we further extend their methods to also consider cubical persistence for rectangular grids and investigate the use of functional summaries for spatial dependence testing as detailed below.
@@ 4.2 One-Sample Permutation Testing for Spatial Dependence
  • $H_{0}$: Expression values are randomly distributed in space.
  • $H_{1}$: Expression values are not spatially random.
We thus retain the original gene expression values as well as the spatial graph and randomly permute which expression value is assigned to each location to simulate spatial randomness, i.e. the lack of any spatial pattern. Then, significance is assessed as follows:
@@
where $\mathds{1}$ is the indicator function summing up the number of times the $L^{p}$ norm, $||\cdot||_{p}$, between the observed Betti curve $\operatorname{\mathrm{b}}_{\text{obs}}$ and the empirical mean Betti curve under the null model $\widebar{\operatorname{\mathrm{b}}_{0}}$ is more extreme than the difference between each individual Betti curve under the null model and their overall mean.
Note that we compute the mean Betti curve given the null hypothesis, $\widebar{\operatorname{\mathrm{b}}_{0}}$, as the mean Betti curve across all $n_{\text{perm}}$ permutations as well as the observed Betti curve, while also adding 1 to both the numerator and denominator. This bias correction is included to avoid zero p-values and bias the test towards more conservative estimates by including this pseudo-count [32]. We thus conduct a two-sided, one-sample hypothesis test for each spatial feature. Whenever multiple genes are compared simultaneously, we further recommend applying a suitable multiple testing adjustment, such as Benjamini-Hochberg procedure to control the false discovery rate [7], which we will apply throughout our experiments. Across our experiments, the number of permutations is set to $1000$, giving a practical compromise with regards to computational efficiency considerations.
Beyond statistical reasoning, there also exist more deterministic approaches to spatially variable gene discovery. Here, the goal is to either rank features by the strength of their spatial dependence or to find a predefined number of the top most spatially variable features. In practice, these approaches are key for both explanatory analysis and for feature selection as a first step to uncover and understand spatial patterns in data [6, 13]. For further analysis, we thus choose to rank spatially expressed features by their empirical p-values, following similar statistical approaches [25].
@@ 4.3 Alternative Spatial Variability Detection Methods
In order to evaluate the performance of our spatial testing approach, we compare our methods with three standard methods for detecting spatial dependence.
@@
Moran’s I is one of the oldest and most widely-used measures for spatial autocorrelation. It is known to be a suitable baseline for SVG benchmarking [25] and can be computed as
$I=\frac{n}{\sum_{i,j}w_{i,j}}\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}w_{i,j}(f_{i}-\bar{f})(f_{j}-\bar{f})}{\sum_{i=1}^{n}(f_{i}-\bar{f})^{2}}.$ (12)
Here, $n$ corresponds to the number of observations, i.e. the number of vertices in the spatial graph $G=(X,E)$. The feature $f_{i}=f(x_{i})$ corresponds to our gene expression measurement for observation $x_{i}\in X$ and $\bar{f}=\frac{1}{n}\sum_{i=1}^{n}f_{i}$ is the mean feature value across observations. We use spatial adjacencies to define the weights $w_{i,j}$, so that $w_{i,j}=1$ if $(x_{i},x_{j})\in E$ and $w_{i,j}=0$ otherwise. The value of Moran’s I is directly interpretable, with $-1$ indicating negative autocorrelation, 00 corresponding to spatial randomness, and $1$ indicating positive autocorrelation. Further, for statistical analysis we use Moran’s I in the same randomised permutation testing framework as the previously introduced topological summaries explained in Section 4.2.
@@
sepal simulates diffusion processes and models a stochastic process that tracks the spread of a feature on the spatial domain over time as it tends towards randomness. Diffusion is simulated on either a rectangular or hexagonal grid and sepal then measures diffusion times, which are used to assess spatial patterns and measure how long it takes the diffusion process to converge. The higher the diffusion time, the stronger the spatial signal. Note that this method does not conduct any statistical testing and thus does not provide p-values. Instead, it provides a ranking of genes by their diffusion times. Moreover, sepal is only applicable to hexagonal or rectangular grids. Hence, whenever data is given with unstructured coordinates, we follow the implementation by Andersson and Lundeberg [6] to convert spatial measurements to a rectangular grid.
@@
Finally, SpatialDE applies spatial covariance testing via Gaussian process regression, a class of models originating from geostatistics. In a nutshell, gene expression is modelled via a multivariate normal model that includes both a non-spatial and a spatial covariance term. The model fitted with spatial covariance is then compared to a null model of spatial independence without this spatial component. Likelihood testing is used and p-values are computed analytically using the $\chi^{2}$-distribution with one degree of freedom. As a multiple testing correction, we follow Svensson et al. [41] and report q-values, which specifically control for the positive false discovery rate [40]. Because SpatialDE assumes normally distributed residual noise, we follow the preprocessing steps by Svensson et al. [41]. First, Anscombe’s transformation, a variance stabilising method for negative binomial data, is applied. Second, log total count values are regressed out to achieve independence from the total counts per spatial location.
@@ 4.4 Datasets and Preprocessing
We use three real transcriptomics datasets for our investigation. First, we analyse spatial data of a mouse olfactory bulb (referred to as MOB, replicate 1) and a breast cancer tumour sample (breast, replicate 1) from Ståhl et al. [39], standard datasets commonly used to test spatial variability detection methods. Both have been sequenced using 1k spatial transcriptomics arrays, which measure spots on rectangular grids. Each spot has a diameter of $100\mu m$ and encompasses 10–100 cells. Further, we study one sample from Kuppe et al. [23] (sample IZP10), which consists of spatial measurements of spots on a hexagonal grid. This sample has been taken from tissues with necrotic areas extracted from the heart of a patient that suffered from a heart attack.
@@
Prior to analysis, preprocessing and quality control of omics data are crucial steps for accounting for technical errors as well as filtering out noise. Indeed, normalisation is a key step in spatial omics data analysis as many downstream tasks, including the detection of spatially variable features, rely on adequate preprocessing choices to account for technical effects [20]. Following best practices [20] and alternative spatial variability studies [16], we decide on the following preprocessing strategy and select all genes that have a minimum total expression count of $10$ and are present in at least 1% of locations. Further, we remove all cells with fewer than $10$ counts. These steps are done to avoid analysing cells or genes for which hardly any biological signals have been detected. Finally, we exclude all ribosomal and mitochondrial genes from analysis as these indicate technical artifacts and are not related to the mRNA, i.e. protein-coding genes that are the focus of transcriptomics data analysis [29]. For each real dataset, we then identify the $500$ most variable genes using $\texttt{seurat}_{\texttt{v3}}$ for further analysis. That is, we select the set of genes that vary most notably across cells [20]. Our goal will then be to detect whether these genes that differ noticeably between dislocated cells also show distinct spatial patterns. Note that for the simulated data we skip these quality control steps as we want to compare all SVG detection methods across all simulated features.
@@
Further, as conventionally done in spatial omics data analysis, we decide to preprocess the expression values, i.e. gene expression count data, in the manner best fitting the SVG detection method [25, 13]. For SpatialDE we apply the normalisation strategy preferred for this method, as detailed in Section 4.3 [41]. For sepal, we apply the recommended shifted logarithm normalisation by applying the transformation $\log(f+2)$ to each feature $f$. This transformation is applied to stabilize the variance across the dataset to counteract variable sampling effects, such as overdispersion, and reduce the skewedness of the feature’s distribution. The pseudo-count of $2$ is added to enable the log-transformation also for zero-values counts. To ensure comparability across spatial variability tests, we similarly choose to compute our persistence-based test as well as Moran’s I from the same $\log(f+2)$-transformed features. Note however, that this preprocessing strategy is not the only viable choice and our method can be flexibly applied across varying preprocessing choices [2]. In general, given that our persistence-based methods directly scale through each gene’s expression values, we highly recommend to apply our framework to transformed and normalised features, but we leave the choice of the best data-dependent transformation to further investigation [2].